Given that
f(x) = (2x-5)/(x+2), g(y)= y^2-3 and p(r)= 10/|r|-1
Find in exact form (a)f(0) (b)p(-4)

Answer by CPhill(1959)

**(a) f(0)**
Substitute x = 0 into the function f(x):
f(0) = (2 * 0 - 5) / (0 + 2)
f(0) = -5 / 2
f(0) = -2.5 or -5/2 (exact form)
**(b) p(-4)**
Substitute r = -4 into the function p(r):
p(-4) = 10 / |-4| - 1
p(-4) = 10 / 4 - 1
p(-4) = 2.5 - 1
p(-4) = 1.5 or 3/2 (exact form)